OpenStudy (anonymous):

Find f of g if f(x) = 2x - 1 and g(x) = x + 2. A. 2x - 3 B. 2x + 1 C. 2x - 1 D. 2x + 3

OpenStudy (anonymous):

\[f(g(x)) = 2g(x) - 1 = 2(x + 2) - 1 = 2x + 3\]

OpenStudy (anonymous):

it is B that's right

OpenStudy (anonymous):

dunn no that can't be. where did I do wrong?